# Review of Demographic Fairness in Face Recognition

### v2

@@ V-C Post-Processing Methods
A typical FR pipeline employs similarity functions to obtain a matching score. Terhörst et al. replaced the conventional similarity function by a fairness-driven neural network classifier [120]. By adding a penalization term in the loss function, their method was able to equalizes score distributions across ethnic groups, reducing intra-ethnic bias while maintaining high recognition performance. In another work [121], Terhörst et al. introduced an unsupervised fair score normalization method based on individual fairness principles, which treats similar individuals similarly. During training, they partitioned the identities in finite number of groups using K-means clustering on face embeddings. At inference, they computed the cluster-specific thresholds for both samples contributing to the score, and these threshold were combined with a global threshold to yield normalized scores. Morales et al. [74] introduced SensitiveNets, a framework that enforces the removal of sensitive attribute information from learned feature representations. By disentangling sensitive attributes from task-relevant features, their method ensures that downstream agents cannot exploit protected information, thereby enhancing both privacy and fairness.
Salvador et al. proposed FairCal: a post-processing method aimed at improving fairness in face verification without retraining or requiring sensitive attribute labels [123]. They clustered image embeddings (obtained from frozen FR network) in an unsupervised manner and applied conditional calibration maps per cluster. They reported reduced false matches while improving overall accuracy. In [124], Conti et al. employed a shallow MLP on top of the frozen FR model as a post-processing strategy to mitigate gender bias in FR. They enhanced the representations of underperforming subgroups using a Fair von Mises-Fisher loss with class-specific variance control. In a recent work, Linghu et al. extended traditional score normalization methods, such as $\mathbb{Z}$ and $\mathbb{T}$ normalization, by incorporating demographic information to enhance fairness in FR systems [125]. Furthermore, they evaluated three cohort-based approaches based on imposter scores, Platt scaling, and bi-modal cumulative distribution functions (CDF). Their findings demonstrated that the proposed method improved fairness across both race and gender demographic groups, particularly at low FMRs.
@@
The works on mitigation are summarized in Table V. For most of the state-of-the-art works aimed towards mitigating accuracy differentials, the FR models have been trained on massive datasets, fine-tuned primarily for accuracy, and are potentially operating near their capacity limits. When a new objective, such as fairness, is introduced, it competes with the existing accuracy objective, often resulting in performance degradation. Some recent methods have demonstrated the potential to improve fairness metrics without sacrificing overall accuracy. These methods, however, often involve architectural modifications, such as the addition of auxiliary modules. It is worth noting that such models, by virtue of their altered structure, differ from their original baselines (with more learnable parameters), and thus may not be directly comparable.
VI Future Directions
@@
Lightweight Models: The lightweight FR models, often used in handheld devices and resource-constrained environments, encounter significant challenges concerning demographic fairness. These systems, crucial for privacy-sensitive applications, often inherit limitations in capacity and architecture, leading to non-equitable performance across demographic groups. Bias in lightweight models has garnered limited research attention, despite their widespread deployment in mobile and IoT devices with varying sensor qualities. Techniques like knowledge distillation (KD) and pruning, while essential for model compression, introduce or amplify bias. For instance, Liu et al. highlighted that KD inherits biases from larger teacher models [128], while pruning has been shown to disproportionately impacts underrepresented groups [129, 130]. Incorporating fairness-aware techniques is crucial for mitigating these issues. Lin et al. [131] introduced FairGRAPE, a pruning method that evaluates network connections with demographic considerations, reducing performance disparities. Caldeira et al. [132] proposed a multi-specialized teacher framework where each teacher model is trained on a specific ethnicity and collectively distills knowledge into a student model. Achieving demographic fairness in lightweight models requires targeted compression strategies and ethical evaluations of demographic-specific impacts. By integrating fairness principles into compression techniques, lightweight FR systems can achieve more equitable outcomes while maintaining efficiency and accuracy.
Quantization of Models and Data: Quantization, a model compression technique, retains the original architecture (as opposed to KD and pruning– which often modify the structure) while reducing parameter precision, producing smaller and faster models. However, it can lead to higher demographic bias by prioritizing global performance over the accurate classification of under-represented groups. Such bias underscores the need for rigorous fairness evaluation across demographic subgroups when deploying compressed models. Quantization converts floating-point (FP) models into lower-precision formats like 8-bit, balancing efficiency and accuracy [133]. It typically consists of two approaches: post-training quantization (PTQ) and quantization-aware training (QAT). Some studies, such as Stoychev et al. demonstrated that 8-bit PTQ maintained fairness and accuracy in gender bias for face expression recognition [133]. However, reducing precision to 6 bits significantly degraded fairness, indicating a trade-off between compression and bias mitigation. Although similar investigations for FR systems remain limited, these findings highlight the need to evaluate and ensure demographic fairness in quantized models.
@@
Low Resolution: The research on demographic fairness in FR has predominantly focused on high-resolution images, often overlooking the challenges posed by low-resolution images typically captured by surveillance cameras or from significant distances. One of the primary impediment to research this issue is the scarcity of low-resolution datasets that include demographic attributes. Consequently, demographic disparities in low-resolution FR remain under-explored, despite their importance in real-world applications. A recent work from Atzori et al. attempted to address this gap by designing a novel framework to investigate demographic bias in low-resolution FR [136]. They trained state-of-the-art FR models on various combinations of high- and low-resolution images. Testing on degraded images from five datasets revealed significant disparities across gender and ethnic groups, underscoring the need for timely interventions in low-resolution FR. It may be noted that their approach involved use of a generative model to convert high-resolution face images into realistic low-resolution counterparts. In [45, 59], the impact of blur and image resolution were investigated in the context of 1:$N$ identification, revealing that increased blurring leads to elevated FPIRs. These studies also noted that the impact of blur varies across demographic attributes, particularly race and gender. The quality degradations of blur and resolution in these studies were simulated using Gaussian blurring and bicubic interpolation, respectively. The importance of low-resolution FR is evident in programs like BRIAR [^3] BRIAR Programme by IARPA., which aim to enhance recognition technologies for challenging scenarios, such as long-distance identification and low-quality image acquisition. To tackle demographic fairness in low-resolution FR, there is a need to develop both datasets and models tailored to these unique use-cases.
@@
Datasets for Training and Evaluation: While demographically balanced datasets alone may not fully eliminate bias, they play a critical role in reducing disparities in FR performance. Achieving fairness requires not only demographic diversity (e.g., across race, gender, and age) but also balance with respect to non-demographic or soft attributes such as hairstyle, facial hair, makeup, and occlusion. Therefore, large-scale datasets that are balanced across both demographic and non-demographic dimensions are essential for developing FR models that are both fair and accurate. However, acquiring such datasets is increasingly challenging due to cost and ethical and privacy concerns surrounding biometric data collection. The commercial sector, in particular, faces difficulties as most available datasets are collected locally from consenting individuals—which are often in limited size and demographic representation. These constraints necessitate innovative approaches to address demographic fairness using smaller or synthetic datasets.
@@
Access to datasets annotated with multiple attributes—both demographic and non-demographic—will prove to be a valuable resource for advancing fairness in biometric recognition. Such datasets enable deeper exploration of the underlying causes of bias by facilitating the analysis of correlations and potential causal relationships among factors that influence recognition performance. Moreover, they provide a foundation for designing effective mitigation strategies. When used for assessment, multi-attribute datasets can help controlling or isolating individual factors, enabling systematic benchmarking of their specific impact on model performance.
@@
Causes of Bias: Recent research indicates that accuracy disparities across demographic groups in FR systems are not solely attributable to inherent physical or biological characteristics. Instead, these disparities are frequently influenced by non-demographic factors– such as hairstyle, cosmetic use, or occlusion patterns– that are shaped by cultural norms, social behaviors, or individual preferences. Despite their relevance, such factors have historically received limited attention, with only a few recent studies beginning to reveal their significant impact on algorithmic performance.
@@
Complex Bias Factors (Intersectionality): The majority of research on mitigation of accuracy differentials in FR, as discussed in Sec V has focused on single demographic attributes, such as race, age, or gender. However, several studies have identified that combination or intersection of various demographic factors causes (or amplifies) bias in FR models (cf. Sec II), while few works, such as [62, 43] have addressed the issue of disparities caused due to multiple or combined attributes. Existing bias mitigation techniques typically target one demographic attribute at a time, achieving measurable improvements in fairness for that specific attribute. However, in most cases, it remains unclear whether such processing inadvertently introduces imbalances in other demographic attributes. For instance, enhancing fairness for gender-related bias may increase disparities linked to ethnicity and vice-versa. This highlights the need for systematic evaluations of the intersectionality of demographic factors, such as race and gender combined. Consequently, developing mitigation methods capable of addressing multiple demographic attributes simultaneously remains an open challenge.
@@
VII Conclusion
In this work, we have systematically explored the issue of demographic fairness or bias in FR through different yet interrelated sections: causes, datasets, assessment metrics, and mitigation strategies. We discussed key contributing factors such as training data imbalance, algorithmic sensitivities, and image quality variations, as well as the growing recognition of non-demographic attributes. Our analysis of datasets emphasized the importance of demographic diversity and the need for multi-attribute annotations to enable deeper understanding of performance disparities. We also reviewed a range of fairness evaluation metrics and mitigation techniques across the FR pipeline.
An important challenge to be noted here is the trade-off between fairness and accuracy. Since most FR models are already optimized for performance, adding fairness objectives often introduces conflicting constraints. Although some architectures sometimes improve both, these gains frequently rely on increased model capacity, rather than an inherent resolution of the fairness-accuracy trade-off. Recent studies have shown that gender-related bias may be explained by soft-biometric attributes rather than inherent biological differences, suggesting that some disparities are socio-cultural in nature. However, such understanding is still limited for attributes like race or skin-tone, largely due to the absence of datasets annotated with both demographic and non-demographic variables. Creating such datasets is essential but resource-intensive; synthetic data may help address this gap, though its effectiveness remains sub-par of real-world data. Improved metrics are also needed to quantify the impact of specific attributes on performance in a consistent and interpretable way. Additionally, most fairness evaluations are conducted on large-scale models, yet future deployments–especially on mobile and edge devices–will require lightweight models that may be more prone to differential treatment. Ensuring fairness under such constraints is critical for real-world adoption.
@@ VII Conclusion
[Dataset URLs] This appendix provides the URLs for the datasets referenced in Section III. These URLs were accessed in March 2025, corresponding to the time of preparation of this woek. It is important to note that many of these datasets are intended solely for non-commercial research purposes. Readers are advised to review the licensing terms prior to use.

### v3

@@ VI-C Post-Processing Methods
A typical FR pipeline employs similarity functions to obtain a matching score. Terhörst et al. replaced the conventional similarity function by a fairness-driven neural network classifier [106]. By adding a penalization term in the loss function, their method was able to equalizes score distributions across ethnic groups, reducing intra-ethnic bias while maintaining high recognition performance. In another work [136], Terhörst et al. introduced an unsupervised fair score normalization method based on individual fairness principles, which treats similar individuals similarly. During training, they partitioned the identities in finite number of groups using K-means clustering on face embeddings. At inference, they computed the cluster-specific thresholds for both samples contributing to the score, and these threshold were combined with a global threshold to yield normalized scores. In a recent work, Linghu et al. extended traditional score normalization methods, such as $\mathbb{Z}$ and $\mathbb{T}$ normalization, by incorporating demographic information to enhance fairness in FR systems [139]. Furthermore, they evaluated three cohort-based approaches based on imposter scores, Platt scaling, and bi-modal cumulative distribution functions (CDF). Their findings demonstrated that the proposed method improved fairness across both race and gender demographic groups, particularly at low FMRs.
@@
The works on mitigation are summarized in Table V. For most of the state-of-the-art works aimed towards mitigating accuracy differentials, the FR models have been trained on massive datasets, fine-tuned primarily for accuracy, and are potentially operating near their capacity limits. When a new objective, such as fairness, is introduced, it competes with the existing accuracy objective, often resulting in performance degradation. This effect is commonly known as the performance-fairness trade-off, where improvements in fairness are often accompanied by reduced recognition accuracy.
Some recent methods have demonstrated the potential to improve fairness metrics without sacrificing overall accuracy. These methods, however, often involve architectural modifications, such as the addition of auxiliary modules. It is important to note that in end-to-end FR pipelines, the total number of model parameters can vary significantly depending on the chosen bias mitigation strategy. Some in-processing methods aim to improve fairness by regularizing or fine-tuning the existing FR network, without adding any additional modules. In contrast, other approaches may integrate substantial additional components, potentially increasing the network depth and capacity. While such designs may yield improved recognition performance along with reduced bias, they differ significantly in architecture and complexity from the baseline models. In addition to structural differences, variations in training data used for both– bias mitigation and FR backbone– play important role. As a result, performance comparisons may not be directly equivalent, and should be made with transparency and careful consideration.
VII Future Directions
@@
Lightweight Models: The lightweight FR models, often used in handheld devices and resource-constrained environments, encounter significant challenges concerning demographic fairness. These systems, crucial for privacy-sensitive applications, often inherit limitations in capacity and architecture, leading to non-equitable performance across demographic groups. Performance disparities in lightweight models has garnered limited research attention, despite their widespread deployment in mobile and IoT devices with varying sensor qualities. Techniques like knowledge distillation (KD) and pruning, while essential for model compression, introduce or amplify bias. For instance, Liu et al. highlighted that KD inherits biases from larger teacher models [142], while pruning has been shown to disproportionately impacts underrepresented groups [143, 144]. Incorporating fairness-aware techniques is crucial for mitigating these issues. Lin et al. [145] introduced FairGRAPE, a pruning method that evaluates network connections with demographic considerations, reducing performance disparities. Caldeira et al. [146] proposed a multi-specialized teacher framework where each teacher model is trained on a specific ethnicity and collectively distills knowledge into a student model. Achieving demographic fairness in lightweight models requires targeted compression strategies and ethical evaluations of demographic-specific impacts. By integrating fairness principles into compression techniques, lightweight FR systems can achieve more equitable outcomes while maintaining efficiency and accuracy.
Quantization of Models and Data: Quantization, a model compression technique, retains the original architecture (as opposed to KD and pruning– which often modify the structure) while reducing parameter precision, producing smaller and faster models. However, it can lead to higher disparities across demographic groups by prioritizing global performance over the accurate classification of under-represented groups. Such performance disparities underscore the need for rigorous fairness evaluation across demographic subgroups when deploying compressed models. Quantization converts floating-point (FP) models into lower-precision formats like 8-bit, balancing efficiency and accuracy [147]. It typically consists of two approaches: post-training quantization (PTQ) and quantization-aware training (QAT). Some studies, such as Stoychev et al. demonstrated that 8-bit PTQ maintained fairness and accuracy in gender bias for face expression recognition [147]. However, reducing precision to 6 bits significantly degraded fairness, indicating a trade-off between compression and bias mitigation. Although similar investigations for FR systems remain limited, these findings highlight the need to evaluate and ensure demographic fairness in quantized models.
@@
Low Resolution: The research on demographic fairness in FR has predominantly focused on high-resolution images, often overlooking the challenges posed by low-resolution images typically captured by surveillance cameras or from significant distances. One of the primary impediment to research this issue is the scarcity of low-resolution datasets that include demographic attributes. Consequently, demographic disparities in low-resolution FR remain under-explored, despite their importance in real-world applications. A recent work from Atzori et al. attempted to address this gap by designing a novel framework to investigate performance differences across demographic groups in low-resolution FR [150]. They trained state-of-the-art FR models on various combinations of high- and low-resolution images. Testing on degraded images from five datasets revealed significant disparities across gender and ethnic groups, underscoring the need for timely interventions in low-resolution FR. It may be noted that their approach involved use of a generative model to convert high-resolution face images into realistic low-resolution counterparts. In [51, 66], the impact of blur and image resolution were investigated in the context of 1:$N$ identification, revealing that increased blurring leads to elevated FPIRs. These studies also noted that the impact of blur varies across demographic attributes, particularly race and gender. The quality degradations of blur and resolution in these studies were simulated using Gaussian blurring and bicubic interpolation, respectively. The importance of low-resolution FR is evident in programs like BRIAR [^3] BRIAR Programme by IARPA., which aim to enhance recognition technologies for challenging scenarios, such as long-distance identification and low-quality image acquisition. To tackle demographic fairness in low-resolution FR, there is a need to develop both datasets and models tailored to these unique use-cases.
@@
Datasets for Training and Evaluation: While demographically balanced training datasets alone may not fully eliminate performance differences in FR, they play a critical role in reducing such disparities. Achieving fairness requires not only demographic diversity (e.g., across race, gender, and age) but also balance with respect to non-demographic or soft attributes such as hairstyle, facial hair, makeup, and occlusion. Therefore, large-scale datasets that are balanced across both demographic and non-demographic dimensions are essential for developing FR models that are both fair and accurate. However, acquiring such datasets is increasingly challenging due to cost and ethical and privacy concerns surrounding biometric data collection. The commercial sector, in particular, faces difficulties as most available datasets are collected locally from consenting individuals—which are often in limited size and demographic representation. These constraints necessitate innovative approaches to address demographic fairness using smaller or synthetic datasets. When it comes to evaluation and benchmarking, many existing datasets are significantly imbalanced– typically over-representing a single demographic group, such as Caucasians. Alongside developing more inclusive training datasets, there is a pressing need to construct evaluation datasets that more accurately reflect demographic diversity.
@@
Access to datasets annotated with multiple attributes–both demographic and non-demographic–will prove to be a valuable resource for advancing fairness in biometric recognition. Such datasets enable deeper exploration of the underlying causes of performance differences across demographic groups by facilitating the analysis of correlations and potential causal relationships among factors that influence recognition performance. Moreover, they provide a foundation for designing effective mitigation strategies. When used for assessment, multi-attribute datasets can help controlling or isolating individual factors, enabling systematic benchmarking of their specific impact on model performance.
@@
Causes of Performance Disparities: Recent research indicates that accuracy disparities across demographic groups in FR systems are not solely attributable to inherent physical or biological characteristics. Instead, these disparities are frequently influenced by non-demographic factors– such as hairstyle, cosmetic use, or occlusion patterns– that are shaped by cultural norms, social behaviors, or individual preferences. Despite their relevance, such factors have historically received limited attention, with only a few recent studies beginning to reveal their significant impact on algorithmic performance.
@@
Complex Bias Factors (Intersectionality): The majority of research on mitigation of accuracy differentials in FR, as discussed in Sec VI has focused on single demographic attributes, such as race, age, or gender. However, several studies have identified that combination or intersection of various demographic factors causes (or amplifies) bias in FR models (cf. Sec III), while few works, such as [69, 52] have addressed the issue of disparities caused due to multiple or combined attributes. Existing bias mitigation techniques typically target one demographic attribute at a time, achieving measurable improvements in fairness for that specific attribute. However, in most cases, it remains unclear whether such processing inadvertently introduces imbalances in other demographic attributes. For instance, enhancing fairness for gender-related bias may increase disparities linked to ethnicity and vice-versa. This highlights the need for systematic evaluations of the intersectionality of demographic factors, such as race and gender combined. Consequently, developing mitigation methods capable of addressing multiple demographic attributes simultaneously remains an open challenge.
@@
VIII Conclusion
In this work, we have systematically explored the issue of performance differences in FR across demographic groups, or bias through different yet interrelated sections: causes, datasets, assessment metrics, and mitigation strategies. We discussed key contributing factors such as training data imbalance, skin-tone variations, and image quality, as well as the growing recognition of non-demographic attributes. Our analysis of datasets emphasized the importance of demographic diversity and the need for multi-attribute annotations to enable deeper understanding of performance disparities. We also reviewed a range of fairness evaluation metrics and mitigation techniques across the FR pipeline.
As highlighted by several recent studies, non-demographic covariates (or soft-biometric factors) can have a significant influence on FR performance, including accuracy and error rates. Given that both training and evaluation datasets, including real-world deployments, are rarely controlled for demographic attributes in isolation, it becomes critical to reconsider how we interpret variations in FR outcomes across demographic groups. Many of these non-demographic factors are closely entangled with demographic attributes (e.g., hairstyles and makeup often vary by culture or gender). Consequently, performance disparities observed across groups may not be attributable to the demographic variable alone. Without precise isolation and identification of contributing factors, attributing lower FR performance to demographic bias may be misleading. Therefore, caution must be exercised before concluding that an FR system exhibits bias toward a particular demographic group.
Another important challenge to be noted here is the trade-off between fairness and accuracy. Since most FR models are already optimized for performance, adding fairness objectives often introduces conflicting constraints. Although some architectures sometimes improve both, these gains frequently rely on increased model capacity, rather than an inherent resolution of the fairness-accuracy trade-off. Recent studies have shown that gender-related performance differences may be explained by soft-biometric attributes rather than inherent biological differences, suggesting that some disparities are socio-cultural in nature. However, such understanding is still limited for attributes like race or skin-tone, largely due to the absence of datasets annotated with both demographic and non-demographic variables. Creating such datasets is essential but resource-intensive; synthetic data may help address this gap, though its effectiveness remains sub-par of real-world data. Improved metrics are also needed to quantify the impact of specific attributes on performance in a consistent and interpretable way. Additionally, most fairness evaluations are conducted on large-scale models, yet future deployments–especially on mobile and edge devices–will require lightweight models that may be more prone to differential treatment. Ensuring fairness under such constraints is critical for real-world adoption.